OpenStudy (anonymous):

Simplify the given expression. square root negative 10 open parentheses 11 plus square root negative 5

OpenStudy (anonymous):

Can you write it down using the equation option please?

OpenStudy (anonymous):

Do you want the explanation or the answer?

OpenStudy (anonymous):

Both please :)

OpenStudy (anonymous):

Okay the answer is the second one.

OpenStudy (anonymous):

and why is that?

OpenStudy (anonymous):

\[(\sqrt{-10})(11+\sqrt{-5}) = 11\sqrt{-10}+(\sqrt{-5})(\sqrt{-10}) = 11\sqrt{-10}+ \sqrt{50} = 11\sqrt{-10} +5\sqrt{2} = (11i \sqrt{10} +5\sqrt{2})\]

OpenStudy (anonymous):

ooo okay thanks i see it :D

OpenStudy (anonymous):

\[11\sqrt{-10} +5\sqrt{2} = 11i \sqrt{10}\]

OpenStudy (anonymous):

+ 5 sqrt(2)

OpenStudy (anonymous):

Yeah i see what you did

OpenStudy (anonymous):

Mind checking over 1 more for me?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Would B, be correct?

OpenStudy (anonymous):

No the correct answer is C.

OpenStudy (anonymous):

O yeah okay yeah because you have to divide the i so it moves top which makes the fraction a negative

OpenStudy (anonymous):

\[(\frac{ 3 }{ 4i }) = (\frac{ 3i }{ 4i^2 }) \] In this case I timed the numerator and denomintor by i

OpenStudy (anonymous):

the i^2 = -1 So it would be -3i/4

OpenStudy (anonymous):

yeah i see. user error forgot i had to do that. Thanks so much !! :D

OpenStudy (anonymous):

No worries. I think a medal would be a good idea though :DDDDD.

OpenStudy (anonymous):

no prob :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Glad to help.